# Learning Satellite Attitude Dynamics with Physics-Informed Normalising Flow

### v2

@@ Abstract
Attitude control is a fundamental aspect of spacecraft operations. Model Predictive Control (MPC) has emerged as a powerful strategy for these tasks, relying on accurate models of the system dynamics to optimize control actions over a prediction horizon. In scenarios where physics models are incomplete, difficult to derive, or computationally expensive, machine learning offers a flexible alternative by learning the system behavior directly from data. However, purely data-driven models often struggle with generalization and stability, especially when applied to inputs outside their training domain. To address these limitations, we investigate the benefits of incorporating Physics-Informed Neural Networks (PINNs) into the learning of spacecraft attitude dynamics, comparing their performance with that of purely data-driven approaches. Using a Real-valued Non-Volume Preserving (Real NVP) neural network architecture with a self-attention mechanism, we trained several models on simulated data generated with the Basilisk simulator. Two training strategies were considered: a purely data-driven baseline and a physics-informed variant to improve robustness and stability. Our results demonstrate that the inclusion of physics-based information significantly enhances the performance in terms of the mean relative error of the best architectures found by 27.08%. These advantages are particularly evident when the learned models are integrated into an MPC framework, where PINN-based models consistently outperform their purely data-driven counterparts in terms of control accuracy and robustness, yielding improvements of up to 42.86% in performance stability error and increased robustness-to-noise.
@@
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
The effectiveness of satellites’ attitude control system (ACS) is critical to the overall result of space missions, impacting both their operational efficiency and life horizon. Accurate pointing capabilities are essential for a wide range of mission objectives, including maintaining reliable communication links, gathering precise data from onboard scientific instruments, and ensuring proper thermal regulation (Wayer et al. (2013)). However, designing robust ACS for satellites remains a significant challenge because of the spacecraft’s complex and non-linear dynamics, as well as the highly variable conditions of the space environment. Among common attitude control actuators, reaction wheels (RWs) are preferred for their high accuracy and moderately fast maneuvers, providing continuous and smooth control (Xia et al. (2017)). However, they also introduce significant nonlinearities, such as frictions and saturation effects. In addition to the complexities introduced by actuators, spacecraft in Earth orbit are continuously subjected to various external disturbance torques, such as gravity gradient, atmospheric drag, and magnetic field torques (Wayer et al. (2013)). These environmental torques increase the angular momentum of the spacecraft, making active control by the ACS essential to maintain the desired orientation. Addressing these complexities has historically relied on two main approaches: model-based methods (Iannelli et al. (2022); Narkiewicz et al. (2020); Pecorilla and Stesina (2024); Romero and de Souza (2019)) and, more recently, data-driven techniques (Cena et al. (2023); Retagne et al. (2024); Wu et al. (2024); Zhang et al. (2025)). However, both exhibit some limitations. Traditional physics-based models, while foundational and offering a structured understanding of the system behavior, often rely on simplifying assumptions. These simplifications can compromise their accuracy and make them difficult to apply effectively in highly dynamic and uncertain environments (Moldabekov et al. (2023)). A notable example of these methods is Model Predictive Control (MPC), which relies on an internal dynamics model, whose fidelity is critical for the stability and robustness of the controller. On the other hand, purely data-driven machine learning (ML) algorithms learn directly from experience without requiring an explicit system model. Although they offer versatility in pattern recognition and scalability, these approaches present significant drawbacks in safety-critical applications. Therefore, improving the spacecraft’s dynamics model is a key enabler for more reliable and effective attitude control, and is the central objective of the approach proposed in this work. In response to the limitations of physics-based models and ML data-driven approaches, Physics-Informed Neural Networks (PINNs) (Raissi et al. (2019)) have emerged as a significant paradigm shift. PINNs integrate the governing physical laws of a system directly into the neural network’s learning process. Traditional PINNs learn to predict the state of the system at a particular time instant, this however does not adapt to control and planning algorithms, which only need to explore the evolution of the actuated system in a limited future time horizon. Therefore alternative approaches have emerged: some merge physical models to AI models trained with data-driven losses (Celestini et al. (2024); Varey et al. (2024)), others add to the data-driven loss an unsupervised physics loss term, which ensures that the physical equations are satisfied at certain points throughout the domain (Antonelo et al. (2024); Falas et al. (2025); Jiang et al. (2024)). This embedding of physical knowledge has a regularization effect, which leads PINN to be more data-efficient and robust.
@@ 1.1 Contribution
    We demonstrate the practical utility of the learned models by embedding them into an MPC framework and evaluating their performance and robustness-to-noise in closed-loop attitude control tasks, significantly limiting the spread of the trajectories when injecting Gaussian noise into the state observations. Moreover, we show that the trained models enable a successful MPC performance even when the spacecraft parameters are subject to estimation errors of up to 10%.
@@ 3.1 Satellite Attitude Dynamics
$S(\omega)=\begin{bmatrix}0&-w_{3}&w_{2}\\ w_{3}&0&-w_{1}\\ -w_{2}&w_{1}&q_{0}\end{bmatrix}$ (1)
@@ 3.2 Problem Formulation
Developing a reliable attitude control system for satellites remains exceptionally challenging given the spacecraft’s inherently nonlinear and intricate dynamic behavior, coupled with the unpredictable conditions of the space environment (Wayer et al. (2013)). Enhancing the accuracy of the spacecraft dynamics model is therefore indispensable for achieving more precise attitude regulation, also adopting advanced model-based control techniques such as model predictive control (MPC).
@@ 3.5 Non-linear MPC with Learned Dynamics
This concrete application has also been carried out with the aim of testing the effect of the learned dynamics in a model-based controller and evaluating its robustness to state estimation errors. Section 4 will describe the metrics used in the two parts of the project in more detail.
@@
$\Delta\hat{\omega}_{t+1}=PINN_{\theta}(X_{t},u_{t},\theta)$ (13)
@@
Where $u_{k}$ is an abbreviation of the commanded RWs torque $u_{rw}$ at time step k, and $\Delta u_{k}$ is the difference between the RWs torque required at the previous time step and the current one. C and R are diagonal cost matrices used to reduce the torques $u_{rw}$ and to limit their variation, smoothing their trajectories. Their non-zero elements are, respectively, equal to 100 and 1, while Q=diag([1000.,1000.,1000.,1000.,1e2,1e2,1e2,1e4,1e4,1e4,1e2,1e2,1e2])Q=diag([1000.,1000.,1000.,1000.,1e-2,1e-2,1e-2,1e-4,1e-4,1e-4,1e-2,1e-2,1e-2]). Finally, n is the horizon of the MPC, which in our experiments is set to $n=10$.
@@ 4 Experimental setups
In this section we describe in depth the dataset used, the NNs architectures and the metrics computed to evaluate the results.
@@ 4.1 Dataset
The rationale for this dataset design is to reproduce a representative operational environment for small satellites performing attitude control solely with RWs. By including both nominal and perturbed configurations, the dataset allows assessing not only nominal performance but also the robustness of the learned models to model uncertainties and off-nominal conditions. We ran 300 runs randomly varying the initial attitude, and orbital position, to generate an initial dataset, and other 50 runs with a different satellite inertia matrix and mass, to test the robustness of the trained models to errors of approximately 10% on all axes, as these parameters are typically not known with absolute precision and may change during the operational life. These trajectories cover a broad portion of the attitude space and reaction wheel state space, providing a diverse training and testing distribution of control inputs and system responses. In each simulation, the satellite started with a random angular velocity of the reaction wheels, sampled uniformly from [-300, 300] rotations per minute. Each simulation lasted 3 minutes with a sampling and control time of 0.1 seconds, and a simulation time step of 0.001 seconds with Runge-Kutta 4 as integrator.
@@ 4.2 Models
4.3 Evaluation Metrics
@@
Secondly, in the experiments named MPC Experiments, we used the best performing neural networks to test various non-linear MPCs evaluated with the following metrics:
  1. 1.
    Absolute Performance Error: the instantaneous value of the performance error at any given time, as defined in ”ECSS-E-ST-60-10C” (European Cooperation for Space Standardization (2008)).
  2. 2.
    Stability Performance Error: the difference between the instantaneous performance error at a given time t and the error value at an earlier time $t-\Delta t$, as defined in ”ECSS-E-HB-60-10A” (European Cooperation for Space Standardization (2010)). This has been computed using $\Delta t=10$.
The performance error is defined as the difference between the desired state and the current one. Finally, we performed a robustness-to-noise experiment (Robustness Experiments) adding a random error of up to 10% from a Gaussian distribution to each state variable. We repeated this 10 times for each model. To evaluate these experiments, we will show the trajectories distribution, along with the averages of the Mean Absolute Error (MAE), final error, and spread.
@@ 4.4 Implementation Details
The experiments were performed on a computer with an Intel i7-9700K 3.60GHz CPU, 64 GB of RAM, and a GeForce RTX 2080 Ti. The models and loss functions have been implemented in torch and we used the ”do-mpc” library to implement the MPC with AI as dynamics estimator. The models have been trained on the GPU using a batch size of 16384, found during the hyper-parameter search, and tested on CPU with an inference time of $0.66\pm 0.03ms$ for the biggest models, i.e. the Real NVP with self-attention mechanism.
@@ 5 Experiments and Results
Here we show and discuss the results obtained in the three groups of experiments. Each section presents the results obtained in the relative experiments using the metrics described in Section 4.
@@ 5.2 MPC Experiments
Given the results presented above, we used the models NVPSA2-LD and NVPSA2-DD as dynamics predictor in a MPC controller to compare the performances of the best performing model trained with and without the physics-informed loss. In Figures 5 and 6 the solid lines represent the resulting trajectories when performing a rest-to-rest maneuver over 360 seconds with the model trained only with the data-driven loss (NVPSA2-DD) and both the data-driven and the physics-informed loss (NVPSA2-LD).
Figure 5: Results of noise introduction with NVPSA2-DD model.
Figure 6: Results of noise introduction with NVPSA2-LD model.
Moreover, Figures 7 and 8 show the absolute performance error for the quaternion, the angular velocity of the satellite, and the velocities of the RWs.
Figure 7: Absolute Performance Error for the model NVPSA2-$D$.
Figure 8: Absolute Performance Error for the model NVPSA2-LD.
Additionally, in Figure 9 we show the distributions of the Stability Performance Errors averaged over the 3 axes for all time steps after the completion of the maneuver. The boxplots were used to reduce the number of figures, while giving an indication of the different levels of stability of the two approaches.
Figure 9: Distribution of the Stability Performance Error average over the 3 axes.
The results show that both models are able to reach the required state in about 50 seconds, but the model trained only with the data-driven loss is less stable, as can be seen both from the solid lines in figure 5 and from Figure 9, where the SPEs have consistently higher variance with respect to those of the physics-informed model NVPSA2-LD, which achieves an average absolute value of the SPE between 16.67% and 42.86% lower.
5.3 Robustness Experiments
Finally, for all models, we ran 10 experiments by adding a random noise of up to 10% to the state variables of the MPC. The results obtained are shown in figures 5 and 6, where the blurred areas represent the dispersion of the trajectories, while the solid lines represent the trajectory obtained without noise.
@@ 5.3 Robustness Experiments
Table 3: Results of the experiments performed to study the robustness-to-noise of the models when used in a MPC with a 10% Gaussian noise in the state. We show the average over 10 runs.
Experiment & Variable MAE Final Error Spread
NVPSA2-DD Attitude 0.0230 0.0090 0.0059
NVPSA2-LD Attitude 0.0191 0.0000 0.0004
NVPSA2-DD Ang. Vel. 0.0023 0.0002 0.0015
NVPSA2-LD Ang. Vel. 0.0021 0.0000 0.0007
NVPSA2-DD RWs Speed 10.2738 0.5826 4.9787
NVPSA2-LD RWs Speed 9.9085 0.1592 2.1158
NVPSA2-DD RWs Torque 0.0021 0.0009 0.0030
NVPSA2-LD RWs Torque 0.0017 0.0000 0.0016
In Table 3 we show the average of the Mean Absolute Error (MAE), of the final error, i.e. the absolute value of the performance error at the last time step, and of the spread of the distributions. The results show that the purely data-driven model NVPSA2-DD is less robust to noise, as can be appreciated by looking at the RWs torques in Figure 5 and at the table 3 in which the physics-informed model consistently outperforms the NVPSA2-DD model.

### v3

# Learning Robust Satellite Attitude Dynamics with Physics-Informed Normalising Flow

@@ Abstract
Attitude control is a fundamental aspect of spacecraft operations. Model Predictive Control (MPC) has emerged as a powerful strategy for these tasks, relying on accurate models of the system dynamics to optimize control actions over a prediction horizon. In scenarios where physics models are incomplete, difficult to derive, or computationally expensive, machine learning offers a flexible alternative by learning the system behavior directly from data. However, purely data-driven models often struggle with generalization and stability, especially when applied to inputs outside their training domain. To address these limitations, we investigate the benefits of incorporating Physics-Informed Neural Networks (PINNs) into the learning of spacecraft attitude dynamics, comparing their performance with that of purely data-driven approaches. Using a Real-valued Non-Volume Preserving (Real NVP) neural network architecture with a self-attention mechanism, we trained several models on simulated data generated with the Basilisk simulator. Two training strategies were considered: a purely data-driven baseline and a physics-informed variant to improve robustness and stability. Our results demonstrate that the inclusion of physics-based information significantly enhances the performance in terms of the mean relative error with the best architectures found by 27.08%. These advantages are particularly evident when the learned models are integrated into an MPC framework, where PINN-based models consistently outperform their purely data-driven counterparts in terms of control accuracy and robustness, and achieve improved settling times when compared to traditional MPC approaches, yielding improvements of up to 62%, when subject to observation noise and RWs friction.
@@
journal: Acta Astronautica
@@ 1 Introduction
The effectiveness of satellites’ attitude control system (ACS) is critical to the overall result of space missions, impacting both their operational efficiency and life horizon. Accurate pointing capabilities are essential for a wide range of mission objectives, including maintaining reliable communication links, gathering precise data from onboard scientific instruments, and ensuring proper thermal regulation (Wayer et al. (2013)). However, designing robust ACS for satellites remains a significant challenge because of the spacecraft’s complex and non-linear dynamics, as well as the highly variable conditions of the space environment. Among common attitude control actuators, reaction wheels (RWs) are preferred for their high accuracy and moderately fast maneuvers, providing continuous and smooth control (Xia et al. (2017)). However, they also introduce significant non-linearities, such as frictions and saturation effects. In addition to the complexities introduced by actuators, spacecraft in Earth orbit are continuously subjected to various external disturbance torques, such as gravity gradient, atmospheric drag, and magnetic field torques (Wayer et al. (2013)). These environmental torques increase the angular momentum of the spacecraft, making active control by the ACS essential to maintain the desired orientation. Addressing these complexities has historically relied on two main approaches: model-based methods (Iannelli et al. (2022); Narkiewicz et al. (2020); Pecorilla and Stesina (2024); Romero and de Souza (2019)) and, more recently, data-driven techniques (Cena et al. (2023); Retagne et al. (2024); Wu et al. (2024); Zhang et al. (2025)). However, both exhibit some limitations. Traditional physics-based models, while foundational and offering a structured understanding of the system behavior, often rely on simplifying assumptions. These simplifications can compromise their accuracy and make them difficult to apply effectively in highly dynamic and uncertain environments (Moldabekov et al. (2023)). A notable example of these methods is Model Predictive Control (MPC), which relies on an internal dynamics model, whose fidelity is critical for the stability and robustness of the controller. On the other hand, purely data-driven machine learning (ML) algorithms learn directly from experience without requiring an explicit system model. Although they offer versatility in pattern recognition and scalability, these approaches present significant drawbacks in safety-critical applications. Therefore, improving the spacecraft’s dynamics model is a key enabler for more reliable and effective attitude control, and is the central objective of the approach proposed in this work. In response to the limitations of physics-based models and ML data-driven approaches, Physics-Informed Neural Networks (PINNs) (Raissi et al. (2019)) have emerged as a significant paradigm shift. PINNs integrate the governing physical laws of a system directly into the neural network’s learning process. Traditional PINNs learn to predict the state of the system at a particular time instant, this however does not adapt to control and planning algorithms, which only need to explore the evolution of the actuated system in a limited future time horizon. Therefore alternative approaches have emerged: some merge physical models to AI models trained with data-driven losses (Celestini et al. (2024); Varey et al. (2024)), others add to the data-driven loss an unsupervised physics loss term, which ensures that the physical equations are satisfied at certain points throughout the domain (Antonelo et al. (2024); Falas et al. (2025); Jiang et al. (2024)). This embedding of physical knowledge has a regularization effect, which leads PINN to be more data-efficient and robust.
@@ 1.1 Contribution
    We demonstrate the practical utility of the learned models by embedding them into an MPC framework and evaluating their performance and robustness-to-noise in closed-loop attitude control tasks, significantly limiting the spread of the trajectories when injecting Gaussian noise into the state observations.
  5. 5.
    We show that the trained models enable improvements of the MPC performance with respect to traditional non-linear and linear dynamics, when the spacecraft is subject to parameters estimation and state observation errors of up to 20% and 3% respectively, and RWs friction.
@@ 3.1 Satellite Attitude Dynamics
$S(\omega)=\begin{bmatrix}0&-w_{3}&w_{2}\\ w_{3}&0&-w_{1}\\ -w_{2}&w_{1}&0\end{bmatrix}$ (1)
@@ 3.2 Problem Formulation
Developing a reliable attitude control system for satellites remains exceptionally challenging given the spacecraft’s inherently non-linear and intricate dynamic behavior, coupled with the unpredictable conditions of the space environment (Wayer et al. (2013)). Enhancing the accuracy of the spacecraft dynamics model is therefore indispensable for achieving more precise attitude regulation, also adopting advanced model-based control techniques such as model predictive control (MPC). This becomes particularly important when the spacecraft state and parameters are subject to errors and estimation inaccuracies.
@@ 3.5 Non-linear MPC with Learned Dynamics
This concrete application has also been carried out with the aim of testing the effect of the learned dynamics in a model-based controller and evaluating its robustness to state estimation errors and model uncertainties. Section 4 will describe the metrics used in the two parts of the project in more detail.
@@
$\Delta\hat{\omega}_{t+1}=PINN_{\theta}(\omega_{t},\omega_{rw,t},\dot{\omega}_{t}\,u_{t},\theta)$ (13)
@@
Where $u_{k}$ is an abbreviation of the commanded RWs torque $u_{rw}$ at time step k, and $\Delta u_{k}$ is the difference between the RWs torque required at the previous time step and the current one. C and R are diagonal cost matrices used to reduce the torques $u_{rw}$ and to limit their variation, smoothing their trajectories. Their non-zero elements are equal to 1e11e-1, while $Q^{13\times 13}=$ diag(10000, 10000, 10000, 10000, $1e^{-2},1e^{-2},1e^{-2},1e^{-4},1e^{-4},1e^{-4},1e^{-2},1e^{-2},1e^{-2}$). Finally, n is the horizon of the MPC, which in our experiments is set to $n=10$.
@@ 4 Experimental setups
In this section we describe in depth the dataset used, the NNs architectures, the traditional MPC implementations, the experiments and the metrics computed to evaluate the results.
@@ 4.1 Dataset
The rationale for this dataset design is to reproduce a representative operational environment for small satellites performing attitude control solely with RWs. By including both nominal and perturbed configurations, the dataset allows assessing not only nominal performance but also the robustness of the learned models to model uncertainties. We ran 300 runs randomly varying the initial attitude, and orbital position, to generate an initial dataset, and other 50 runs with a different satellite inertia matrix and mass, to test the robustness of the trained models to errors of approximately 10% on all axes, as these parameters are typically not known with absolute precision and may change during the operational life. These trajectories cover a broad portion of the attitude space and reaction wheel state space, providing a diverse training and testing distribution of control inputs and system responses. In each simulation, the satellite started with a random angular velocity of the reaction wheels, sampled uniformly from [-300, 300] rotations per minute. Each simulation lasted 3 minutes with a sampling and control time of 0.1 seconds, and a simulation time step of 0.001 seconds with Runge-Kutta 4 as integrator.
@@ 4.2 Models
4.3 Traditional MPC
In this work we compare our framework to two different MPC without AI, one uses a state space form (linear MPC) and another one uses a non-linear approach. Given the MPC context introduced in Section 3.5, we use the same input state and cost matrices with the non-linear dynamics, implemented following Section 3.1, and change them when using the state space form. In particular, we used as input state $x=\{q,w\}$ and all terms of matrix $Q^{6\times 6}$ were set to 1000, all other parameters were left unchanged. To implement the discrete state space model, given in eq. 18, we followed (Pecorilla and Stesina (2024)).
$\dot{x}_{t+1}=A_{d}x_{t}+B_{d}u_{t}$ (18)
In eq. 18 matrix $A_{d}$ and matrix $B_{d}$ are computed from eq. 19 and eq. 20 using the zero-order hold control implementation, i.e. keeping the control torque constant according to the control frequency, as shown in eq. 22. Given $w_{s}$, which is the scalar angular velocity of the spacecraft about the Earth’s center, and I, which is the satellite inertia matrix $I_{s}$.
$A=\begin{bmatrix}\begin{array}[]{cc}0^{3\times 3}&0.5I^{3\times 3}\end{array}\\ M^{3\times 6}\\ \end{bmatrix}$ (19)
$B=\begin{bmatrix}0^{3\times 3}\\ \text{diag}\left(\frac{1}{I_{0}},\frac{1}{I_{1}},\frac{1}{I_{2}}\right)\end{bmatrix}$ (20)
where M is defined in eq. 21.
$M=\begin{bmatrix}-8w_{s}^{2}\frac{I_{1}-I_{2}}{I_{0}}&0&0&0&0&w_{s}\frac{I_{0}+I_{2}-I_{1}}{I_{0}}\\ 0&-6w_{s}^{2}\frac{I_{0}-I_{2}}{I_{1}}&0&0&0&0\\ 0&0&-2w_{s}^{2}\frac{I_{1}-I_{0}}{I_{2}}&-w_{s}\frac{I_{2}+I_{0}-I_{1}}{I_{2}}&0&0\end{bmatrix}$ (21)
$A_{d}=e^{A\delta t},B_{d}=\int_{0}^{\delta t}e^{A\sigma}d\sigma$ (22)
4.4 Evaluation Metrics & Experiments
@@
Secondly, in the experiments named MPC Experiments, we used the best performing neural networks to test various non-linear MPCs evaluated with the Stability Performance Error, which is the difference between the instantaneous performance error at a given time t and the error value at an earlier time $t-\Delta t$, as defined in ”ECSS-E-HB-60-10A” (European Cooperation for Space Standardization (2010)). This has been computed using $\Delta t=10$. The performance error is defined as the difference between the desired state and the current one.
We performed 300 Monte Carlo simulations for each model to test their robustness-to-noise, adding a random error of up to 3% from a Gaussian distribution to each state variable, and a continuously distributed random error of up to 10% to the inertia matrix, and up to 20% to the satellite mass. Moreover, friction was inserted in the RWs following a linear dynamics up to $50\%\pm 12,5\%$ of the RWs maximum velocity. The initial attitude was varied randomly in the range ($\frac{\pi}{8}$, $\frac{\pi}{2}$). It should be noted that the neural networks were trained with data generated in simulations without noise and RWs friction. To evaluate these experiments, we will show the trajectories distribution and the stability performance error, along with the average steady-state error over the last minute of simulation and the time required to converge to a attitude error below 1 degree (Settling Time).
@@ 4.5 Implementation Details
The experiments were performed on a computer with an Intel i7-9700K 3.60GHz CPU, 64 GB of RAM, and a GeForce RTX 2080 Ti. The models and loss functions have been implemented in torch and we used the ”do-mpc” library to implement the MPC with AI as dynamics estimator. The models have been trained on the GPU using a batch size of 16384, found during the hyper-parameter search, and tested on CPU with an inference time of $0.66\pm 0.03$ms for the biggest models, i.e. the Real NVP with self-attention mechanism. Finally, the models were tested on relevant edge devices to acquire data about their computational cost. These results are shown in Section 5.3.
@@ 5 Experiments and Results
Here we show and discuss the results obtained in the two groups of experiments. Each section presents the results obtained in the relative experiments using the metrics described in Section 4.
@@ 5.2 MPC Experiments
Given the results presented above, we used the models NVPSA2-LD and NVPSA2-DD as dynamics predictor in a MPC controller to compare the performances of the best performing model trained with and without the physics-informed loss. Comparing them with the two traditional MPC controllers introduced in Section 4.3. In Figure 5 we present the trajectories of 300 Monte Carlo simulations when performing rest-to-rest maneuvers over 360 seconds with the model trained only with the data-driven loss (NVPSA2-DD), both the data-driven and the physics-informed loss (NVPSA2-LD), and the two traditional MPC controllers. In these simulations we randomly changed the orbital position, the initial satellite attitude, as well as introduce parameter estimation errors for the satellite weight (up to 20%) and for the satellite inertia matrix (up to 10%). Moreover, state estimation was subject to a Guassian random noise of up to 3% and the reaction wheels were subject to friction, as described in Section 4.4.
(a) NVPSA2-$D$
(b) NVPSA2-LD
(c) Linear MPC
(d) Non-Linear MPC
Figure 5: 300 MC simulation with parameter estimation errors, state estimation noise and RWs friction for NVPSA2-$D$ (upper left), NVPSA2-LD (upper right), traditional MPC with linear (lower left) and non-linear (lower right) dynamics.
@@
Table 3: Results of the experiments performed to study the robustness-to-noise of the models when used in a MPC with several sources of noise. We show the median over 300 MC simulations.
Experiment
Steady-State Error
[degrees]
Settling Time [s]
NVPSA2-$D$ 0.5865 65.0
NVPSA2-LD 0.0464 42.8
Non-Linear MPC 0.1215 186.2
Linear MPC 0.0024 114.1
The results show that all models are able to reach the required state with a steady-state error below 1 degree, but the model trained only with the data-driven loss is less stable. Additionally, it can be seen that the MPC with the physics-informed model (NVPSA2-LD) is able to reach a good convergence error, given the noise present in the environment, in the minimum amount of time among all controller tested. Though, it tends to produce higher torques even in steady-state. In Table 3 we show the median settling time, computed as the time required to reach a pointing error below 1 degree, and the median of the steady-state error over the last minute of the trajectory, i.e. the angle in degrees between the current and target quaternion. The results show that the MPC with state space model requires considerably more time to reach convergence, while being able to reach a more accurate attitude, with respect to the MPC with NVPSA2-LD, which is the fastest, i.e. 114 seconds for the state space model against the 43 seconds required by NVPSA2-LD. Moreover, the median settling time for the traditional non-linear MPC is approximately 3 minutes. Finally, Figure 6 shows the stability performance error for the quaternion, the angular velocity of the satellite, the velocities of the RWs and torque applied to the RWs when using the physics-informed model NVPSA2-LD.
Figure 6: Stability Performance Error for the MPC with NVPSA2-LD under parameter estimation inaccuracies, RWs friction and state estimation noise.
5.3 Edge Inference
In Table 4 we show the mean and standard deviation of the inference latency for the three neural network architectures in 3 edge devices: Jetson Nano, Raspberry PI 4B and AMD GX-412HC. In the Jetson Nano we tested 4 different combinations of maximum power consumption and target hardware, i.e. CPU or GPU. Given the small size of the model, the results suggest that it is better to perform inference only in the CPU, because the overhead associated to the data movement to the GPU is higher than the improvement in performance. When considering only the CPUs it can be seen that the bigger model (NVPSA2) could be used with a frequency between 100Hz and 380Hz.
Table 4: Inference latency tests on different edge devices and architectures. For each device and architecture, we show the average and standard deviation of the latency, in milliseconds, over 1000 inferences.
Device RealNVP[ms] NVPSA1[ms] NVPSA2[ms]
[…]